Arctic Routing Chip

Computation Structures Group Memo 373 March 7, 1994

G. Andrew Boughton

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.